\begin{thebibliography}{42}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Airola et~al.(2019)Airola, Pohjankukka, Torppa, Middleton,
  Nyk{\"a}nen, Heikkonen, and Pahikkala]{airola2019spatial}
Antti Airola, Jonne Pohjankukka, Johanna Torppa, Maarit Middleton, Vesa
  Nyk{\"a}nen, Jukka Heikkonen, and Tapio Pahikkala.
\newblock The spatial leave-pair-out cross-validation method for reliable {AUC}
  estimation of spatial classifiers.
\newblock \emph{Data Mining and Knowledge Discovery}, 33\penalty0 (3):\penalty0
  730--747, 2019.

\bibitem[Boschetti et~al.(2016)Boschetti, Stehman, and
  Roy]{boschetti2016stratified}
Luigi Boschetti, Stephen~V Stehman, and David~P Roy.
\newblock A stratified random sampling design in space and time for regional to
  global scale burned area product validation.
\newblock \emph{Remote sensing of environment}, 186:\penalty0 465--478, 2016.

\bibitem[Brenning(2022)]{brenning2022spatial}
Alexander Brenning.
\newblock Spatial machine-learning model diagnostics: A model-agnostic
  distance-based approach.
\newblock \emph{International Journal of Geographical Information Science},
  pp.\  1--23, 2022.

\bibitem[Brus(2021)]{brus2021statistical}
Dick~J Brus.
\newblock Statistical approaches for spatial sample survey: Persistent
  misconceptions and new developments.
\newblock \emph{European Journal of Soil Science}, 72\penalty0 (2):\penalty0
  686--703, 2021.

\bibitem[Burke et~al.(2021)Burke, Driscoll, Lobell, and Ermon]{burke2021using}
Marshall Burke, Anne Driscoll, David~B Lobell, and Stefano Ermon.
\newblock Using satellite imagery to understand and promote sustainable
  development.
\newblock \emph{Science}, 371\penalty0 (6535):\penalty0 eabe8628, 2021.

\bibitem[Chi et~al.(2022)Chi, Fang, Chatterjee, and
  Blumenstock]{chi2022microestimates}
Guanghua Chi, Han Fang, Sourav Chatterjee, and Joshua~E Blumenstock.
\newblock Microestimates of wealth for all low-and middle-income countries.
\newblock \emph{Proceedings of the National Academy of Sciences}, 119\penalty0
  (3):\penalty0 e2113658119, 2022.

\bibitem[Dowman \& Reuter(2017)Dowman and Reuter]{dowman2017global}
Ian Dowman and Hannes~I Reuter.
\newblock Global geospatial data from earth observation: Status and issues.
\newblock \emph{International Journal of Digital Earth}, 10\penalty0
  (4):\penalty0 328--341, 2017.

\bibitem[Draidi~Areed et~al.(2022)Draidi~Areed, Price, Arnett, and
  Mengersen]{draidi2022spatial}
Wala Draidi~Areed, Aiden Price, Kathryn Arnett, and Kerrie Mengersen.
\newblock Spatial statistical machine learning models to assess the
  relationship between development vulnerabilities and educational factors in
  children in queensland, australia.
\newblock \emph{BMC Public Health}, 22\penalty0 (1):\penalty0 1--12, 2022.

\bibitem[Fourcade et~al.(2018)Fourcade, Besnard, and
  Secondi]{fourcade2018paintings}
Yoan Fourcade, Aur{\'e}lien~G Besnard, and Jean Secondi.
\newblock Paintings predict the distribution of species, or the challenge of
  selecting environmental predictors and evaluation statistics.
\newblock \emph{Global Ecology and Biogeography}, 27\penalty0 (2):\penalty0
  245--256, 2018.

\bibitem[Gaetan \& Guyon(2010)Gaetan and Guyon]{gaetan2010spatial}
Carlo Gaetan and Xavier Guyon.
\newblock \emph{Spatial statistics and modeling}, volume~90.
\newblock Springer, 2010.

\bibitem[Haining(2009)]{haining2009special}
Robert Haining.
\newblock The special nature of spatial data.
\newblock \emph{The SAGE Handbook of Spatial Analysis. Los Angeles: SAGE
  Publications}, pp.\  5--23, 2009.

\bibitem[Hoffimann et~al.(2021)Hoffimann, Zortea, De~Carvalho, and
  Zadrozny]{hoffimann2021geostatistical}
J{\'u}lio Hoffimann, Maciel Zortea, Breno De~Carvalho, and Bianca Zadrozny.
\newblock Geostatistical learning: Challenges and opportunities.
\newblock \emph{Frontiers in Applied Mathematics and Statistics}, 7:\penalty0
  689393, 2021.

\bibitem[Jean et~al.(2016)Jean, Burke, Xie, Davis, Lobell, and
  Ermon]{jean2016combining}
Neal Jean, Marshall Burke, Michael Xie, W~Matthew Davis, David~B Lobell, and
  Stefano Ermon.
\newblock Combining satellite imagery and machine learning to predict poverty.
\newblock \emph{Science}, 353\penalty0 (6301):\penalty0 790--794, 2016.

\bibitem[Klemmer \& Neill(2021)Klemmer and Neill]{klemmer2021auxiliary}
Konstantin Klemmer and Daniel~B Neill.
\newblock Auxiliary-task learning for geographic data with autoregressive
  embeddings.
\newblock In \emph{Proceedings of the 29th International Conference on Advances
  in Geographic Information Systems}, pp.\  141--144, 2021.

\bibitem[Koh et~al.(2021)Koh, Sagawa, Marklund, Xie, Zhang, Balsubramani, Hu,
  Yasunaga, Phillips, Gao, et~al.]{koh2021wilds}
Pang~Wei Koh, Shiori Sagawa, Henrik Marklund, Sang~Michael Xie, Marvin Zhang,
  Akshay Balsubramani, Weihua Hu, Michihiro Yasunaga, Richard~Lanas Phillips,
  Irena Gao, et~al.
\newblock Wilds: A benchmark of in-the-wild distribution shifts.
\newblock In \emph{International Conference on Machine Learning}, pp.\
  5637--5664. PMLR, 2021.

\bibitem[Le~Rest et~al.(2014)Le~Rest, Pinaud, Monestiez, Chadoeuf, and
  Bretagnolle]{le2014spatial}
K{\'e}vin Le~Rest, David Pinaud, Pascal Monestiez, Jo{\"e}l Chadoeuf, and
  Vincent Bretagnolle.
\newblock Spatial leave-one-out cross-validation for variable selection in the
  presence of spatial autocorrelation.
\newblock \emph{Global ecology and biogeography}, 23\penalty0 (7):\penalty0
  811--820, 2014.

\bibitem[Lohr(2021)]{lohr2021sampling}
Sharon~L Lohr.
\newblock \emph{Sampling: design and analysis}.
\newblock Chapman and Hall/CRC, 2021.

\bibitem[Maxwell et~al.(2021)Maxwell, Warner, and
  Guill{\'e}n]{maxwell2021accuracy}
Aaron~E Maxwell, Timothy~A Warner, and Luis~Andr{\'e}s Guill{\'e}n.
\newblock Accuracy assessment in convolutional neural network-based deep
  learning remote sensing studies—part 1: Literature review.
\newblock \emph{Remote Sensing}, 13\penalty0 (13):\penalty0 2450, 2021.

\bibitem[Meyer \& Pebesma(2022)Meyer and Pebesma]{meyer2022machine}
Hanna Meyer and Edzer Pebesma.
\newblock Machine learning-based global maps of ecological variables and the
  challenge of assessing them.
\newblock \emph{Nature Communications}, 13\penalty0 (1):\penalty0 1--4, 2022.

\bibitem[Meyer et~al.(2019)Meyer, Reudenbach, W{\"o}llauer, and
  Nauss]{meyer2019importance}
Hanna Meyer, Christoph Reudenbach, Stephan W{\"o}llauer, and Thomas Nauss.
\newblock Importance of spatial predictor variable selection in machine
  learning applications--moving from data reproduction to spatial prediction.
\newblock \emph{Ecological Modelling}, 411:\penalty0 108815, 2019.

\bibitem[Mil{\`a} et~al.(2022)Mil{\`a}, Mateu, Pebesma, and
  Meyer]{mila2022nearest}
Carles Mil{\`a}, Jorge Mateu, Edzer Pebesma, and Hanna Meyer.
\newblock Nearest neighbour distance matching leave-one-out cross-validation
  for map validation.
\newblock \emph{Methods in Ecology and Evolution}, 2022.

\bibitem[Mitchell et~al.(2019)Mitchell, Wu, Zaldivar, Barnes, Vasserman,
  Hutchinson, Spitzer, Raji, and Gebru]{mitchell2019model}
Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman,
  Ben Hutchinson, Elena Spitzer, Inioluwa~Deborah Raji, and Timnit Gebru.
\newblock Model cards for model reporting.
\newblock In \emph{Proceedings of the conference on fairness, accountability,
  and transparency}, pp.\  220--229, 2019.

\bibitem[Nakalembe(2018)]{nakalembe2018characterizing}
Catherine Nakalembe.
\newblock Characterizing agricultural drought in the {K}aramoja subregion of
  {U}ganda with meteorological and satellite-based indices.
\newblock \emph{Natural Hazards}, 91\penalty0 (3):\penalty0 837--862, 2018.

\bibitem[Nikparvar \& Thill(2021)Nikparvar and Thill]{nikparvar2021machine}
Behnam Nikparvar and Jean-Claude Thill.
\newblock Machine learning of spatial data.
\newblock \emph{ISPRS International Journal of Geo-Information}, 10\penalty0
  (9):\penalty0 600, 2021.

\bibitem[Nilsen et~al.(2021)Nilsen, Tejedor-Garavito, Leasure, Utazi,
  Ruktanonchai, Wigley, Dooley, Matthews, and Tatem]{nilsen2021review}
Kristine Nilsen, Natalia Tejedor-Garavito, Douglas~R Leasure, C~Edson Utazi,
  Corrine~W Ruktanonchai, Adelle~S Wigley, Claire~A Dooley, Zoe Matthews, and
  Andrew~J Tatem.
\newblock A review of geospatial methods for population estimation and their
  use in constructing reproductive, maternal, newborn, child and adolescent
  health service indicators.
\newblock \emph{BMC health services research}, 21\penalty0 (1):\penalty0 1--10,
  2021.

\bibitem[Oliver et~al.(2021)Oliver, Meyer, Ranipeta, Winner, and
  Jetz]{oliver2021global}
Ruth~Y Oliver, Carsten Meyer, Ajay Ranipeta, Kevin Winner, and Walter Jetz.
\newblock Global and national trends, gaps, and opportunities in documenting
  and monitoring species distributions.
\newblock \emph{PLoS Biology}, 19\penalty0 (8):\penalty0 e3001336, 2021.

\bibitem[Paliyam et~al.(2021)Paliyam, Nakalembe, Liu, Nyiawung, and
  Kerner]{paliyam2021street2sat}
Madhava Paliyam, Catherine Nakalembe, Kevin Liu, Richard Nyiawung, and Hannah
  Kerner.
\newblock Street2sat: A machine learning pipeline for generating ground-truth
  geo-referenced labeled datasets from street-level images.
\newblock In \emph{Tackling Climate Change with Machine Learning Workshop at
  the International Conference on Machine Learning}, 2021.

\bibitem[Ploton et~al.(2020)Ploton, Mortier, R{\'e}jou-M{\'e}chain, Barbier,
  Picard, Rossi, Dormann, Cornu, Viennois, Bayol, et~al.]{ploton2020spatial}
Pierre Ploton, Fr{\'e}d{\'e}ric Mortier, Maxime R{\'e}jou-M{\'e}chain, Nicolas
  Barbier, Nicolas Picard, Vivien Rossi, Carsten Dormann, Guillaume Cornu,
  Ga{\"e}lle Viennois, Nicolas Bayol, et~al.
\newblock Spatial validation reveals poor predictive performance of large-scale
  ecological mapping models.
\newblock \emph{Nature communications}, 11\penalty0 (1):\penalty0 1--11, 2020.

\bibitem[Pohjankukka et~al.(2017)Pohjankukka, Pahikkala, Nevalainen, and
  Heikkonen]{pohjankukka2017estimating}
Jonne Pohjankukka, Tapio Pahikkala, Paavo Nevalainen, and Jukka Heikkonen.
\newblock Estimating the prediction performance of spatial models via spatial
  k-fold cross validation.
\newblock \emph{International Journal of Geographical Information Science},
  31\penalty0 (10):\penalty0 2001--2019, 2017.

\bibitem[Proctor et~al.(2023)Proctor, Carleton, and Sum]{NBERw30861}
Jonathan Proctor, Tamma Carleton, and Sandy Sum.
\newblock Parameter recovery using remotely sensed variables.
\newblock Working Paper 30861, National Bureau of Economic Research, January
  2023.
\newblock URL \url{http://www.nber.org/papers/w30861}.

\bibitem[Roberts et~al.(2017)Roberts, Bahn, Ciuti, Boyce, Elith,
  Guillera-Arroita, Hauenstein, Lahoz-Monfort, Schr{\"o}der, Thuiller,
  et~al.]{roberts2017cross}
David~R Roberts, Volker Bahn, Simone Ciuti, Mark~S Boyce, Jane Elith, Gurutzeta
  Guillera-Arroita, Severin Hauenstein, Jos{\'e}~J Lahoz-Monfort, Boris
  Schr{\"o}der, Wilfried Thuiller, et~al.
\newblock Cross-validation strategies for data with temporal, spatial,
  hierarchical, or phylogenetic structure.
\newblock \emph{Ecography}, 40\penalty0 (8):\penalty0 913--929, 2017.

\bibitem[Robinson et~al.(2022)Robinson, Ortiz, Park, Lozano, Kaw, Sederholm,
  Dodhia, and Ferres]{robinson2022fast}
Caleb Robinson, Anthony Ortiz, Hogeun Park, Nancy Lozano, Jon~Kher Kaw, Tina
  Sederholm, Rahul Dodhia, and Juan M~Lavista Ferres.
\newblock Fast building segmentation from satellite imagery and few local
  labels.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  1463--1471, 2022.

\bibitem[Rolf et~al.(2020)Rolf, Jordan, and Recht]{rolf2020post}
Esther Rolf, Michael~I Jordan, and Benjamin Recht.
\newblock Post-estimation smoothing: A simple baseline for learning with side
  information.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, pp.\  1759--1769. PMLR, 2020.

\bibitem[Rolf et~al.(2021)Rolf, Proctor, Carleton, Bolliger, Shankar, Ishihara,
  Recht, and Hsiang]{rolf2021generalizable}
Esther Rolf, Jonathan Proctor, Tamma Carleton, Ian Bolliger, Vaishaal Shankar,
  Miyabi Ishihara, Benjamin Recht, and Solomon Hsiang.
\newblock A generalizable and accessible approach to machine learning with
  global satellite imagery.
\newblock \emph{Nature communications}, 12\penalty0 (1):\penalty0 1--11, 2021.

\bibitem[Schratz et~al.(2019)Schratz, Muenchow, Iturritxa, Richter, and
  Brenning]{schratz2019hyperparameter}
Patrick Schratz, Jannes Muenchow, Eugenia Iturritxa, Jakob Richter, and
  Alexander Brenning.
\newblock Hyperparameter tuning and performance assessment of statistical and
  machine-learning algorithms using spatial data.
\newblock \emph{Ecological Modelling}, 406:\penalty0 109--120, 2019.

\bibitem[Sofaer et~al.(2019)Sofaer, Jarnevich, Pearse, Smyth, Auer, Cook,
  Edwards~Jr, Guala, Howard, Morisette, et~al.]{sofaer2019development}
Helen~R Sofaer, Catherine~S Jarnevich, Ian~S Pearse, Regan~L Smyth, Stephanie
  Auer, Gericke~L Cook, Thomas~C Edwards~Jr, Gerald~F Guala, Timothy~G Howard,
  Jeffrey~T Morisette, et~al.
\newblock Development and delivery of species distribution models to inform
  decision-making.
\newblock \emph{BioScience}, 69\penalty0 (7):\penalty0 544--557, 2019.

\bibitem[Song \& Kim(2022)Song and Kim]{song2022three}
Insang Song and Daehyun Kim.
\newblock Three common machine learning algorithms neither enhance prediction
  accuracy nor reduce spatial autocorrelation in residuals: An analysis of
  twenty-five socioeconomic data sets.
\newblock \emph{Geographical Analysis}, 2022.

\bibitem[Stehman et~al.(2021)Stehman, Pengra, Horton, and
  Wellington]{stehman2021validation}
Stephen~V Stehman, Bruce~W Pengra, Josephine~A Horton, and Danika~F Wellington.
\newblock Validation of the us geological survey's land change monitoring,
  assessment and projection (lcmap) collection 1.0 annual land cover products
  1985--2017.
\newblock \emph{Remote Sensing of Environment}, 265:\penalty0 112646, 2021.

\bibitem[Tuia et~al.(2016)Tuia, Persello, and Bruzzone]{tuia2016domain}
Devis Tuia, Claudio Persello, and Lorenzo Bruzzone.
\newblock Domain adaptation for the classification of remote sensing data: An
  overview of recent advances.
\newblock \emph{IEEE geoscience and remote sensing magazine}, 4\penalty0
  (2):\penalty0 41--57, 2016.

\bibitem[Valavi et~al.(2018)Valavi, Elith, Lahoz-Monfort, and
  Guillera-Arroita]{valavi2018blockcv}
Roozbeh Valavi, Jane Elith, Jos{\'e}~J Lahoz-Monfort, and Gurutzeta
  Guillera-Arroita.
\newblock blockcv: An r package for generating spatially or environmentally
  separated folds for k-fold cross-validation of species distribution models.
\newblock \emph{bioRxiv}, pp.\  357798, 2018.

\bibitem[Wadoux et~al.(2021)Wadoux, Heuvelink, De~Bruin, and
  Brus]{wadoux2021spatial}
Alexandre MJ-C Wadoux, Gerard~BM Heuvelink, Sytze De~Bruin, and Dick~J Brus.
\newblock Spatial cross-validation is not the right way to evaluate map
  accuracy.
\newblock \emph{Ecological Modelling}, 457:\penalty0 109692, 2021.

\bibitem[Yuan \& McKee(2022)Yuan and McKee]{yuan2022embedding}
May Yuan and Arlo McKee.
\newblock Embedding scale: New thinking of scale in machine learning and
  geographic representation.
\newblock \emph{Journal of Geographical Systems}, 24\penalty0 (3):\penalty0
  501--524, 2022.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2023_mlrs_workshop_mine,times}
\pdfoutput=1

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{amsmath}
\usepackage{cleveref}

\title{Evaluation challenges for geospatial ML}

\author{Esther Rolf \\%
Harvard University
}

\newcommand{\loc}[0]{\ell}
\newcommand{\evalset}[0]{\mathcal{S}_{\textrm{eval}}}
\newcommand{\pop}[0]{\mathcal{D}}

\iclrfinalcopy 
\begin{document}

\maketitle

\begin{abstract}
As geospatial machine learning models and maps derived from their predictions are increasingly used for downstream analyses in science and policy, it is imperative to evaluate their accuracy and applicability. 
Geospatial machine learning has key distinctions from other learning paradigms, and as such, the correct way to measure performance of spatial machine learning outputs has been a topic of debate. 
In this paper, I delineate unique challenges of model evaluation for geospatial machine learning with global or remotely sensed datasets, 
culminating
in concrete takeaways to improve  evaluations of geospatial model performance. 
%
%

\end{abstract}

\section{Motivation}
\label{sec:intro}

Geospatial machine learning (ML), for example with remotely sensed data, is being used across consequential domains, including  
public health \citep{nilsen2021review,draidi2022spatial}
conservation  \citep{sofaer2019development}, food security \citep{nakalembe2018characterizing}, and wealth estimation 
\citep{jean2016combining,chi2022microestimates}. 
By both their use and their very nature, geospatial predictions have a purpose beyond model benchmarking;
mapped data are to be read, scrutinized, and acted upon.
Thus, it is critical to rigorously and comprehensively evaluate how well a predicted map represents the state of the world it is meant to reflect, or how well a spatial ML model performs across the many conditions in 
which it might be used.
%

Unique structures in remotely sensed and geospatial data complicate or even invalidate use of traditional ML evaluation procedures. 
Partially as a result of misunderstandings of these complications, 
the stated performance of several geospatial models and predictive maps has come into question \citep{
fourcade2018paintings, ploton2020spatial}.
This in turn has sparked disagreement on what the ``right'' evaluation procedure  is. 
With respect to a certain set of spatial evaluation methods (described in \S\ref{sec:spatial_crossval}), one is  jointly presented with the arguments that ``spatial cross-validation is essential in preventing overoptimistic model performance''  \citep{meyer2019importance} and ``spatial cross-validation methods have
no theoretical underpinning and should not be used for assessing map accuracy'' \citep{wadoux2021spatial}. 
That both statements can simultaneously hold reflects the importance of using a diverse set of evaluation methods tailored to the many ways in which a geospatial ML model might be used.
%

In this paper, I situate the challenges of geopsatial model evaluation in the perspective of an ML researcher, synthesizing prior work across ecology, geology, statistics, and machine learning. 
I aim in part to disentangle key factors that complicate effective evaluation of model and map performance.
First and foremost, evaluation procedures should be designed to measure as closely as possible the quantity or phenomena they are intended to assess (\S\ref{sec:map_accuracy_vs_model_performance}).
After the relevant performance measures are established, considerations can be made about 
what is feasible with the available data (\S\ref{sec:data_section}).
%
With all of this in mind, possible evaluation procedures (\S\ref{sec:spatial_eval_methods}) can be compared and tailored to the task at hand.
Recognizing the interaction of these  distinct but related steps  exposes opportunities to improve geospatial performance assessment, both in individual studies and more broadly (\S\ref{sec:discussion_oppurtuntities}).

\section{Map accuracy and model performance: Contrasting views}
\label{sec:map_accuracy_vs_model_performance} 
Estimating accuracy 
indices and  corresponding  
uncertainties 
of geospatial predictions is essential to reporting geospatial ML performance (\S\ref{subsec:map_accuracy}), especially when prediction maps will be used for downstream analyses 
or policy decisions.
At the same time, the potential value of a geospatial ML model likely extends beyond that of a single mapped output (\S\ref{sec:prediction_regimes}).
Delineating the (possibly many) facets of desired model and map use is key to measuring geospatial ML performance
 (\S\ref{sec:contrasting_views}).


\subsection{Map accuracy as a population parameter to be estimated}
\label{subsec:map_accuracy}
%
%

%
Establishing notation we will use throughout, let $\hat{y}(\loc)$ denote a model's predicted value  at location $\loc$, and $y(\loc)$ the reference, or ``ground truth'' value (which we assume can be measured). 
To calculate a \textbf{map accuracy index as a population parameter} 
for accuracy index $F$ is to calculate
$
A(\pop) = F(\{ (\hat{y}(\loc), y(\loc))\}_{\loc \in \pop})
$
where $\pop$ is the \textbf{target population} of map use (e.g. all (lat, lon) pairs in a global grid, or all administrative units in a set of countries). 
Examples of common $F$ 
include root mean squared error,
and
area under the ROC curve, among many others \citep{maxwell2021accuracy}. 

Typically, one only has a limited set of values $y$ for locations in an evaluation set $\loc \in \evalset$ from which to compute a statistic
$ \hat{A}(\evalset) $
to estimate $A(\pop)$.
%
%
%
%
\citet{wadoux2021spatial} discuss the value of using a design-independent probability sample for design-based estimation of $A$ (in contrast, model-based estimation makes statistical  assumptions about the data \citep{brus2021statistical}).
Here a \textbf{design-independent sample} is one collected independently of the model training process. 
A \textbf{probability sample} is one for which every location in $\pop$ has 
a positive probability of appearing in $\evalset$, and these probabilities are known for all $\ell \in \evalset$ 
%
(see, e.g. \citet{lohr2021sampling}).
\citet{wadoux2021spatial} emphasize that 
when $\evalset$ is a design independent probability sample from population $\pop$,
design-based inference can be used to estimate $A(\mathcal{\pop})$ with $\hat{A}(\evalset)$, 
\emph{regardless of the prediction model or distribution of training data}.

Computing statistically valid estimates of map accuracy indices is clearly a key component of reporting overall geospatial ML model performance.
It is often important 
to understand how accuracy and uncertainty in predictions vary across sub-populations $\pop_{r_1},  \pop_{r_2} \ldots \subset \pop$ (such as administrative regions or climate zones \citep{meyer2022machine}).
If \textbf{local accuracy} indexes $A(\pop_{r_1}),  A(\pop_{r_2}) \ldots$ are low in certain sub-regions, this could expose concerns about fairness or model applicability. 
%

\subsection{Model performance extends beyond map accuracy}
\label{sec:prediction_regimes}
Increasingly, geospatial ML models are designed with the goal of being used \emph{outside} of the regions where training labels are available. Models trained with globally available remotely sensed data might be used to ``fill in'' spatial gaps common to other data modalities (\S\ref{sec:data_availability}). The goals of \textbf{spatial generalization}, \textbf{spatial extrapolation} or \textbf{spatial domain adaption} can take different forms: e.g. applying a model trained with data  from one region 
to a wholly new region, 
or using data from a few clusters or subregions to extend predictions across the entire region. 
When spatial generalizability is desired, performance should be assessed specifically with respect to this goal (\S\ref{sec:spatial_eval_methods}).

%

%

While spatial generalization is a key component of performance for many geospatial models,
it too is just one facet of geospatial model performance.
Proposed uses of geospatial ML models and their outputs include estimation of natural or causal parameters \citep{NBERw30861}, and reducing autocorrelation of prediction residuals in-sample \citep{song2022three}. 
Other important facets of geospatial ML performance  are model interpretability \citep{brenning2022spatial} and usability, including the resources required to train, deploy and maintain models \citep{rolf2021generalizable}.  
%

%

\subsection{Contrasting perspectives on performance assessment}
\label{sec:contrasting_views}
The differences between estimating map accuracy as a population parameter (\S\ref{subsec:map_accuracy})
and assessing a model's performance in the conditions it is most likely to be used 
(\S\ref{sec:prediction_regimes})
are central to one of the
discrepancies introduced in \S\ref{sec:intro}.
%
%
\citet{meyer2019importance, ploton2020spatial, meyer2022machine} state concerns in light of numerous ecological studies applying non-spatial validation techniques with the explicit purpose of spatial generalization. 
They rightly caution that when data exhibit spatial correlation (\S\ref{sec:spatial_data_structures}), non-spatial validation methods will almost certainly over-estimate predictive performance in these use cases.
\citet{wadoux2021spatial}, in turn,
argue that performance metrics from spatial validation methods will
not necessarily tell you anything about $A$ as a population parameter.

A second discrepancy between these two perspectives hinges on what data is assumed to be available (or collectable).
%
While there are some major instances of probability samples being collected for evaluation of global-scale maps  \citep{boschetti2016stratified,stehman2021validation}, this is far from standard standard  in geospatial ML studies \citep{maxwell2021accuracy}. 
More often, datasets are created ``by merging all data available from different sources'' \citep{meyer2022machine}.
Whatever the intended use of a geospatial model, the availability of and structures within geopsatial and remotely sensed data must be contended with in order to reliably evaluate any sort of performance.

\section{Structures and patterns in spatial and remotely sensed data}
\label{sec:data_section}
Geospatial and remotely sensed data exhibit distinct structures. 
For example, the chosen extent and scale of a spatial prediction unit
 ($\ell$ in \S\ref{sec:map_accuracy_vs_model_performance}) has important implications for the design, use, and evaluation of geospatial ML models, evidenced by the phenomena of  ``modifiable areal unit problem'' and ``ecological fallacy'' \citep{haining2009special,nikparvar2021machine,yuan2022embedding}.
Here, I focus on two key factors of geospatial data that affect 
the validity of
geospatial ML evaluation methods.

\subsection{Spatial structures: (Auto)correlation and covariate shift}
\label{sec:spatial_data_structures}
One key phenomena exhibited by many geopsatial data is that values of a variable (e.g. tree canopy height) are often correlated across locations. Formally, for random process $Z$, the  \textbf{spatial autocorrelation function} is defined as 
$
R_{ZZ}(\loc_i, \loc_j) = \mathbb{E}[Z(\loc_i) Z(\loc_j)]  / \sigma_i \sigma_j 
$, where $\sigma_i, \sigma_j$ are the standard deviations associated with $Z(\loc_i),Z(\loc_j)$.
%
%
%
For geospatial variables, we might expect $R_{ZZ}(\loc_i, \loc_j) > 0$ when $\loc_i$ and $\loc_j$ are closer together, namely that values of $Z$ at nearby points tend to be closer in value.
%
The degree of spatial autocorrelation in data 
can be assessed with statistics such as Moran's $I$ and  Geary's $C$, and semi-variogram analsyes (see, e.g. \citep{gaetan2010spatial}). 

Spatial autocorrelations and  correlations between predictor and label variables can be an important source of structure to leverage in geospatial ML models \citep{rolf2020post,klemmer2021auxiliary}, yet they also present challenges.
%
Models can ``over-rely'' on spatial correlations in the data, leading to over-estimated  accuracy despite poor spatial generalization performance. 
%
Overfitting to spatial relationships in training data is of particular concern in the when data distributions differ between training regions and regions of use. 
Such \textbf{covariate shifts} are common in geospatial data, e.g. across climate zones, or spectral shifts in satellite imagery  \citep{tuia2016domain,hoffimann2021geostatistical}.

Presence of spatial correlations or domain shift alone do not invalidate  assessing map accuracy with a probability sample  (\S\ref{subsec:map_accuracy}). 
%
%
However, when evaluation is limited to existing data, issues of data availability and representivity can amplify the challenges of geospatial model evaluation.

\subsection{Availability, quality, and representivity of geospatial evaluation data}
\label{sec:data_availability}

Many geospatial datasets exhibit gaps in coverage or quality of data.
\citet{meyer2022machine} evidence trends of geographic clustering around research sites primarily in a small number of countries, across
three datasets used for global mapping in ecology. 
%
%
\citet{oliver2021global} find geographical bias in coverage of species distribution data  aggregated from 
field observation, sensors measurements, and citizen science efforts. 
\citet{burke2021using}
note that the frequency at which nationally representative data on agriculture, population, and economic factors are collected varies widely across the globe.  
%
%
While earth observation data such as satellite imagery have comparatively much higher coverage across time and space \citep{burke2021using}, coverage of commercial products has been shown to be biased toward regions of high commercial value \citep{dowman2017global}.


Filling in data gaps is goal for which  geospatial ML can be transformative (\S\ref{sec:prediction_regimes}), yet these same gaps complicate model evaluation. 
When training data are clustered in small regions, this can affect our ability to train a high-performing model. 
When \emph{evaluation data} are clustered in small regions, this affects our ability to evaluate geospatial ML model performance at all. 




\section{Spatially-aware evaluation methods: A brief overview}

In \S\ref{sec:data_section}, we established that geospatial data  generally exhibit spatial correlations and 
data gaps, even when target use areas $\pop$ are small.
It is well documented that 
calculating accuracy indices with non-spatial validation methods (e.g. standard k-fold cross-validation) will generally \emph{over-estimate}  performance in such settings. 
%
%
%
%
Spatially-aware evaluation methods can control the spatial distribution of training and validation set points to better simulate conditions of intended model use.

\label{sec:spatial_eval_methods}
\subsection{Spatial cross-validation methods}
\label{sec:spatial_crossval}

Several spatial cross-validation methods have been proposed 
that reduce spatial dependencies between
train set points $\ell \in \mathcal{S}_{\textrm{train}}$ from evaluation set points $\ell \in \mathcal{S}_{\textrm{eval}}$.
%
Spatial cross-validation methods
typically stratify training and evaluation instances by larger geographies \citep{roberts2017cross, valavi2018blockcv}
e.g. existing boundaries, spatial blocks, or automatically generated clusters.
Buffered cross-validation methods (such as spatial leave-one-out \citep{le2014spatial}, leave-pair out \citep{airola2019spatial} and k-fold cross validation \citep{pohjankukka2017estimating}) control the minimum distance from any training point to any evaluation point.
%
In addition to evaluating model performance, 
spatial cross-validation has also been suggested as a way to to improve model selection and parameter estimation in geospatial ML \citep{meyer2019importance,schratz2019hyperparameter, roberts2017cross}.

While separating $\ell \in \mathcal{S}_{\textrm{train}}$ 
from $\ell \in \mathcal{S}_{\textrm{eval}}$ can  reduce the amount of correlation between training and evaluation data, a spatial split also induces a higher degree of spatial extrapolation to the learning setup and potentially reduces variation in the evaluation set labels.
%
As a result, it is possible for spatial validation methods to systematically \emph{under-report} performance, especially in interpolation regimes \citep{roberts2017cross, wadoux2021spatial}. 
%
In a different flavor from the evaluation methods above, \citet{mila2022nearest} propose to match the distribution of nearest neighbor distances between train and evaluation set points to the corresponding distances between train set and target use area. 


\subsection{Other spatially-aware validation evaluation methods}
\label{subsec:other_spatial_eval}
%
%
When the intended use of a geospatial model is to generate predictions outside the training distribution, it is critical to test the model's ability to generalize across different conditions.
For example, studies have varied the amount of spatial extrapolation required by changing parameters of the spatial validation setups in \S\ref{sec:spatial_crossval} , e.g. with buffered leave one out \citep{ploton2020spatial,brenning2022spatial} and checkerboard designs \citep{roberts2017cross,rolf2021generalizable}. 
\citet{jean2016combining} assess extrapolation ability across pairs of countries by iteratitvely training in one region and evaluating performance in another. \citet{rolf2021generalizable} find that the distances at which a geospatial model has extrapolation power can differ substantially depending on the properties of the prediction variable.
%

%

It is always critical to put the reported performance of geospatial ML models in context.
Visualizing the spatial distributions of predictions and error residuals can help expose overreliance on spatially correlated predictors \citep{meyer2019importance} and sub-regions with low local model performance.
Comparing performance to that of a baseline model built entirely on spatial predictors can contextualize the value-add of a new geospatial model \citep{fourcade2018paintings, rolf2021generalizable}.
%
%

\section{Taking stock: Considerations and opportunities}
\label{sec:discussion_oppurtuntities}
Comprehensive reporting of performance is critical for geospatial ML methods, 
especially as stated gains in research progress make their way to maps and decisions of real-world consequence. 
Evaluating performance  of geospatial models is especially challenging in the face of spatial correlations and limited availability or representivity of data. 
This means non-spatial data splits are generally unsuitable for geospatial model evaluation with most existing datasets. 
Spatially-aware validation methods are an important indicator of model performance including spatial generalization; 
however, 
they generally do not provide valid statistical estimates of prediction map accuracy. 
%
%
%
This brings us to end with three key opportunities for improving the landscape of geospatial ML evaluation: 

\textbf{Opportunity 1: Invest in evaluation data} to measure map accuracy and overall performance of geospatial models.
When remote annotations are appropriate, labeling tools (e.g. \citet{robinson2022fast}) can facilitate the creation of probability-sampled evaluation datasets.
Data collection and aggregation efforts can focus on filling existing geospatial data gaps \citep{paliyam2021street2sat} or simulating real-world prediction conditions like covariate or domain shift \citep{koh2021wilds}.

\textbf{Opportunity 2: Invest in evaluation frameworks} to precisely and transparently and report performance and valid uses of a geospatial ML model (\`{a} la 
``model cards'' \citep{mitchell2019model}).
This includes improving spatial baselines, expanding methods for reporting
uncertainty over space, and delineating ``areas of applicability" for geospatial models, e.g. as in \citet{meyer2022machine}.

%

\textbf{Opportunity 3:} If the available data and evaluation frameworks are insufficient, \textbf{explain the limitations of what types of performance can be evaluated}. Distinguish between performance measures that estimate a statistical parameter and those that indicate potential skill for a possible use case. 

\subsubsection*{Acknowledgments}
Esther Rolf was supported by the Harvard Data Science Initiative (HDSI) and the Center for Research
on Computation and Society (CRCS). Thank you to Konstantin Klemmer,  Caleb Robinson, and Jessie Finocchiaro for feedback on earlier drafts of this work.
%
%
\bibliography{spatial_references}
\bibliographystyle{iclr2023_conference}

\end{document}